OpenStudy (mmend98):

Which of the following is not true about beta emissions? Difference of the atomic number by one between the old element and new element formed Are written as 0-1e Cannot penetrate concrete Less penetrating than alpha emissions

OpenStudy (mmend98):

@Preetha

OpenStudy (mmend98):

@taramgrant0543664

OpenStudy (taramgrant0543664):

Well which one are you thinking it is? I'll cross one off for you, beta emissions are unable to pass through concrete but it can pass through something like paper.

OpenStudy (taramgrant0543664):

So it is not C

OpenStudy (mmend98):

Its asking what it not is. So C would be the answer right?

OpenStudy (zale101):

Here's a hint: Beta particles are high speed electrons, travels faster, and are less heavily ionized than alpha particles.

OpenStudy (taramgrant0543664):

C says cannot penetrate concrete so that's true the question asks what is not true

OpenStudy (mmend98):

oh lol sorry

OpenStudy (mmend98):

Thanks Zale it would be D

OpenStudy (taramgrant0543664):

It's ok!

OpenStudy (taramgrant0543664):

And yes it would be D

OpenStudy (zale101):

Yes it is D. because since beta particles are less heavily ionized then that makes it more penetrating than alpha particles. Alpha particles are thickly ionized and are less penetrating.

OpenStudy (mmend98):

Given the following chemical equation, how many moles of carbon are necessary to react with 5.52 moles of SO2? 5C(s) + 2SO2 (g) arrow CS2 (l) + 4CO (g) 5.52 moles SO2 2.00 moles SO2 13.8 moles SO2 11.0 moles SO2 this one would be 13.8 right ? because the formula is 5.52molesSO2×5molesC2molesSO2=13.8

OpenStudy (zale101):

are you familiar wit mole to mole ratio?

OpenStudy (zale101):

with*

OpenStudy (mmend98):

yea im familiar did i do it right ? thats what im asking lol

OpenStudy (zale101):

Yes. Sorry, i did not see it.

OpenStudy (mmend98):

thats ok mate

OpenStudy (mmend98):

Fusion refers to the splitting of an atomic nucleus into fragments of roughly the same mass amount. True False

OpenStudy (taramgrant0543664):

13.8 would be correct

OpenStudy (mmend98):

Fusion refers to the splitting of an atomic nucleus into fragments of roughly the same mass amount. True False

OpenStudy (zale101):

What do you think? True or false? Make any educated guess

OpenStudy (zale101):

Search Nuclear Fission and see if it matches the definition.

OpenStudy (mmend98):

the educated guess i would make is false

OpenStudy (mmend98):

i just never know with this weird wording tbh

OpenStudy (taramgrant0543664):

That is referring to fission which is the splitting, fusion is the colliding of atoms

OpenStudy (zale101):

Yep. False because what Tara stated above ^

OpenStudy (mmend98):

Given a solution with a pH of 2.8, what is the [H+] of the solution? 2.8 M 6.31 x 10-12 M 1.58 x 10-3 M 11.2 M

OpenStudy (mmend98):

this one would be c?

OpenStudy (taramgrant0543664):

Yes it would be c [H+]=10^-2.8 Which gives c

OpenStudy (mmend98):

A solution is highly concentrated if there is: a large volume of solution a large amount of solvent and a small amount of solute a large amount of solute and a small amount of solvent a high density value for the solution

OpenStudy (mmend98):

im confused on this one for some reason tara lol

OpenStudy (mmend98):

youve been unreal help btw

OpenStudy (taramgrant0543664):

Well when you think about it volume of solution doesn't matter because it could be at any volume and density value doesn't make sense for it either. Think about the difference between something dilute and something that is concentrated and that will sort out between B and C

OpenStudy (mmend98):

it would be b

OpenStudy (taramgrant0543664):

And no problem I enjoy helping and once my SS gets a little higher I am planning on applying for the qualified helper in chemistry so I'm glad to hear that I'm helping!

OpenStudy (mmend98):

it would be b right ?

OpenStudy (taramgrant0543664):

dilute: a solution that contains a small proportion of solute relative to solvent concentrated: a solution that contains a large proportion of solute relative to solvent.

OpenStudy (mmend98):

oh lol its c

OpenStudy (mmend98):

im dumb af

OpenStudy (taramgrant0543664):

No you're not it was a little mistake that happens, I've made a lot of mistakes it happens but that's how you learn

OpenStudy (mmend98):

Calculate the change in freezing point when 116.8 grams of barium chloride are added to 250 grams of water. Kf = 1.86°C/m; Kb = 0.512°C/m -4.17°C -12.5°C -8.35°C 0°C last question im stuck on

OpenStudy (taramgrant0543664):

ΔT= Kf · m · i So ΔT is the change in temperature that you solve for Kf is given m is the molarity to find this you need to find the moles of barium chloride and you need to convert 250g of water into kg. then you do moles divided by kg of water to find m i is the van't Hoff dissociation factor so BaCl2 dissociates into Ba2+ +2Cl- let me double check this i, I can't quite remember the rule but the rest is right I'll look it up quick

OpenStudy (taramgrant0543664):

So BaCl2 breaks into Ba2+ and 2Cl- so it breaks into 3 parts so i is 3 but NaCl for example breaks into Na+ and Cl- so i=2

OpenStudy (taramgrant0543664):

If we put it all together ΔT= Kf · m · i ΔT= (1.86)m(3) so m just needs to be found

OpenStudy (mmend98):

i got A ?

OpenStudy (taramgrant0543664):

How did you get a negative? What did you get for m?

OpenStudy (mmend98):

wait i got D i forgot to multiply one term.

OpenStudy (taramgrant0543664):

Oh wait it's suppose to be negative don't mind me it's a little late not really thinking for a second there lol

OpenStudy (mmend98):

then i think it has to be a.\

OpenStudy (taramgrant0543664):

I got something different what did you get for m?

OpenStudy (mmend98):

what did u get then ill check mine over

OpenStudy (taramgrant0543664):

I got 2.24 as my m value

OpenStudy (mmend98):

i got 5.2

OpenStudy (mmend98):

idek how

OpenStudy (taramgrant0543664):

116.8g/ molar mass of BaCl2 (208.233)=0.5609 0.5609/ kg of water so 250/1000 this =2.24

OpenStudy (mmend98):

i got B

OpenStudy (mmend98):

yes your correct

OpenStudy (taramgrant0543664):

Yes B would be correct!

OpenStudy (mmend98):

Given a solution with a pOH of 13.6, what is the [OH-] of the solution? 13.6 M 2.51 x 10-14 M 0.4 M 3.98 x 10-1 M would be a?

OpenStudy (mmend98):

im sorry if its too late go ahead and have a great night !! u were amazing

OpenStudy (taramgrant0543664):

I can check this last one quick for you just give me one sec

OpenStudy (mmend98):

thanks Tara

OpenStudy (taramgrant0543664):

pOH=-log[OH-] 13.6=-log[OH-] -13.6=log[OH-] 10^(-13.6)=[OH-] It's not A

OpenStudy (mmend98):

B is what i got

OpenStudy (mmend98):

see this is why i ask u ;) haha

OpenStudy (taramgrant0543664):

Yes B is correct

OpenStudy (mmend98):

so its 2.51 x 10-14 M ?

OpenStudy (mmend98):

sweet

OpenStudy (taramgrant0543664):

Yes it is!